# Large Language Models for Crash Detection in Video: A Survey of Methods, Datasets, and Challenges

### v1

@@ I Introduction
Crash detection is vital for autonomous vehicles, traffic monitoring, and emergency response. Traditionally, it has relied on heuristics or vision-based CNNs using motion cues. However, these methods struggle with generalization, occlusion, and complex temporal dependencies.
Large Language Models (LLMs), such as GPT-4 and PaLM, have demonstrated impressive reasoning across modalities, especially when integrated with visual encoders like CLIP, Flamingo, or BLIP. This integration introduces a new paradigm: interpreting crash-related events in video via language-guided reasoning, captioning, and event prediction.
This survey aims to provide a structured overview of recent approaches leveraging LLMs/VLMs for crash understanding in video, summarizing models, datasets, and evaluation strategies.
@@ II Background and Related Work
The domain of video-based crash detection has evolved significantly, moving from classical computer vision techniques to modern deep learning and language-guided approaches. This section provides background on the foundational methods that preceded LLM integration and reviews the key developments in Vision-Language Models (VLMs) and their application to video understanding, which together set the stage for the methods surveyed in this paper.
II-A Crash Detection in Video
Classical approaches to crash detection in video include frame differencing, optical flow, background subtraction, and CNN-based classification. These methods rely heavily on low-level motion or scene cues, such as changes in pixel intensity, motion vectors, or spatial patterns, to identify anomalies indicative of crashes. Early research in this area often focused on trajectory-based analysis, where anomalies were detected if objects deviated from learned normal paths [1, 2]. However, these methods were sensitive to occlusion and relied on continuous tracking.
Subsequent approaches shifted to using low-level features that were more robust. Frame differencing detects sudden changes between consecutive frames, often used to identify abrupt vehicle movements [3]. Optical flow analyzes motion patterns to track vehicle trajectories, enabling detection of collisions or erratic behavior [4, 5]. Background subtraction isolates moving objects from static scenes, useful for detecting crash-related anomalies in traffic videos [6]. Other traditional methods leveraged hand-crafted features like Histograms of Oriented Gradients (HOG) and Histograms of Optical Flow (HOF) to capture appearance and motion information [7].
While effective in controlled settings, these approaches often struggle with the ambiguity and variability of real-world driving scenarios, such as varying lighting conditions or complex traffic interactions [8, 9]. This led to the adoption of deep learning, with CNN-based classification, such as using ResNet or YOLO, leveraging deep learning to classify video frames or regions as crash or non-crash events, often combining spatial and temporal features through architectures like Convolutional Long Short-Term Memory (Conv-LSTM) [3, 10]. However, even these methods can be limited, prompting the exploration of more robust methods like Large Language Models (LLMs) and Vision-Language Models (VLMs) for enhanced crash detection [11].
@@ II-A Crash Detection in Video
Domain fine-tunesCrashLM-24
(fine-tuned GPT-4V)
@@
II-B Vision-Language Models
Vision-Language Models (VLMs) integrate image or video encoders with text decoders or bidirectional transformers to enable multimodal understanding, combining visual and textual data for tasks such as video understanding and anomaly detection. Early VLMs, such as CLIP [23] and ALIGN [24], focus on matching images with text descriptions, leveraging large-scale image-text pairs to learn robust cross-modal representations. These models excel in tasks like image classification and retrieval but lack temporal reasoning capabilities essential for video-based applications. Newer systems, including BLIP-2 [25], Flamingo [26], and Video-LLaMA [27], incorporate spatiotemporal reasoning, making them well-suited for crash detection in video data. These advancements allow VLMs to process sequential video frames, capture dynamic patterns, and reason about complex traffic scenarios, addressing the limitations of classical approaches like frame differencing and CNN-based classification [3, 11].
II-B1 Early VLMs: Image-Text Alignment
Early VLMs, such as CLIP and ALIGN, rely on contrastive learning to align image and text embeddings in a shared feature space. CLIP, for instance, is trained on 400 million image-text pairs, enabling zero-shot classification by matching visual content to textual descriptions [23]. In the context of crash detection, CLIP can classify individual video frames as crash or non-crash by comparing frame embeddings to textual prompts like “car accident” or “normal driving” [28]. Similarly, ALIGN uses a larger dataset of 1.8 billion image-text pairs to achieve robust alignment, improving performance on diverse visual scenarios [24]. However, these models are limited to static images, requiring additional temporal processing for video-based crash detection.
II-B2 Advanced VLMs: Spatiotemporal Reasoning
Recent VLMs, such as BLIP-2, Flamingo, and Video-LLaMA, extend beyond static image processing to incorporate spatiotemporal reasoning, critical for analyzing dynamic crash events in videos. BLIP-2 leverages a bootstrapped approach, combining pre-trained vision and language models to enhance multimodal tasks like video question answering and action recognition [25]. Its ability to process video frames as sequences enables detection of temporal anomalies, such as sudden vehicle collisions. Flamingo, built on a mixture of vision and language datasets, uses a gated cross-attention mechanism to integrate visual and textual inputs, achieving strong performance on video understanding tasks [26]. Video-LLaMA, specifically designed for video, employs a video encoder to capture temporal dynamics, aligning video features with language representations for tasks like crash detection [27]. These models can process video sequences to identify crash-related patterns, such as abrupt stops or collisions, by combining visual cues with textual reasoning.
II-B3 Application to Crash Detection
VLMs enhance crash detection by integrating visual and textual data to reason about complex traffic scenarios. For example, Video-LLaMA can analyze a sequence of dashcam frames to detect anomalies, using textual prompts to guide the identification of crash events [27]. Similarly, frameworks like VERA use VLMs such as LLaVA-1.5 to verbalize video content, generating explanations for detected anomalies (e.g., “a car collided with another vehicle”) [29]. More recent approaches leverage VLMs to extract rich semantic features from video frames, which are then used to identify subtle anomalies that traditional methods might miss. For instance, some methods use VLMs to generate textual descriptions of scenes, which are then analyzed for anomalous keywords or phrases [9]. These approaches outperform classical methods, which struggle with variability in lighting or traffic conditions [6, 10]. By leveraging spatiotemporal reasoning and multimodal integration, VLMs provide a robust framework for real-time crash detection in intelligent transportation systems.
II-C Large Language Models for Video Understanding

### v2

@@ Abstract
LatencyAccuracy100500100015002000+80859095Deployment ReadyConditionalResearch OnlyCrashLLMVERAVideo-LLaMALAVADHolmes-VADCRASHIDEAL ZONE
Figure 2: Accuracy vs. Latency Trade-off Analysis for LLM-based Crash Detection Methods. The green zone represents deployment-ready systems (<500<500 ms), yellow shows conditional deployment scenarios, and red indicates research-only systems.
@@ I Introduction
Crash detection from video data has been a critical challenge in intelligent transportation systems for over two decades. Early approaches, dating back to the early 2000s, relied on classical computer vision techniques such as optical flow analysis, background subtraction, and trajectory-based anomaly detection [1, 2, 3]. These methods, while computationally efficient and interpretable, achieved limited accuracy (70-85%) and struggled with fundamental challenges including environmental sensitivity, occlusion handling, and poor generalization across different traffic scenarios [4, 5].
The deep learning revolution (2015-2020) brought significant improvements, with CNN-based approaches and specialized architectures like Two-Stream networks, 3D CNNs, and SlowFast networks achieving 85-95% accuracy on video analysis benchmarks [6, 7, 8, 9]. Video anomaly detection methods, particularly those developed for surveillance applications, provided direct foundations for crash detection through weakly-supervised learning and future frame prediction approaches [10, 11, 12]. However, these methods remained limited by their black-box nature, substantial data requirements, and inability to provide interpretable explanations for safety-critical decisions.
The emergence of multimodal learning (2020-2022) began to address some of these limitations through vision-language models that could incorporate semantic understanding alongside visual processing [13, 14, 15, 16]. These approaches demonstrated the potential for cross-modal reasoning and zero-shot generalization, though they remained constrained by limited temporal modeling capabilities and basic language understanding.
Large Language Models (LLMs), such as GPT-4 and PaLM, have recently demonstrated impressive reasoning capabilities across modalities, especially when integrated with visual encoders like CLIP, Flamingo, or BLIP [17, 18, 19]. This integration introduces a transformative paradigm: interpreting crash-related events in video through language-guided reasoning, detailed captioning, causal analysis, and predictive event modeling. Unlike previous approaches that focused primarily on detection accuracy, LLM-based systems can provide rich explanations (e.g., “A vehicle swerved due to sudden braking, causing a rear-end collision”), perform causal reasoning about crash sequences, and adapt to new incident types through few-shot learning.
However, as Table I illustrates, this evolution has come with new challenges. While LLM-based approaches offer unprecedented capabilities in contextual understanding and explanation generation, they face significant deployment barriers including high computational costs (200ms-2s latency), hallucination risks, and substantial memory requirements that limit real-world applicability.
This survey provides the first comprehensive analysis of this fundamental paradigm shift, examining how the field has evolved from reactive, pixel-based anomaly detection to proactive, context-aware event interpretation powered by language-guided reasoning. We present a unified taxonomy of LLM-based approaches, analyze their performance against classical and deep learning baselines, and critically assess their readiness for real-world deployment. Our analysis reveals both the transformative potential and the significant engineering challenges that must be addressed to realize practical LLM-based crash detection systems.
TABLE I: Evolution of Crash Detection Methods: Performance and Limitations Across Different Eras
Era Period Typical Accuracy Latency Key Methods Main Advantages Critical Limitations Classical 2000-2015 70-85% <<100ms Optical flow, background subtraction, trajectory analysis, SVM Fast processing, interpretable rules, low computational cost Manual feature engineering, environmental sensitivity, poor generalization Deep Learning 2015-2020 85-95% 100-500ms CNN, 3D CNN, Two-Stream, I3D, SlowFast, LSTM Automatic feature learning, better generalization, robust to variations Black-box nature, data hungry, limited context understanding Early Multimodal 2020-2022 88-92% 300-800ms CLIP, UNITER, VideoCLIP, BLIP Cross-modal understanding, zero-shot capabilities, semantic reasoning Limited temporal modeling, basic language understanding LLM-based 2023-2025 85-92% 200ms-2s Video-LLaMA, VERA, CrashLLM, Holmes-VAD, HybridMamba Rich explanations, causal reasoning, contextual understanding, few-shot learning High computational cost, hallucination risk, deployment challenges
@@ II Background and Related Work
Video-based crash detection has evolved through distinct phases: from classical computer vision techniques (2000-2015) to deep learning approaches (2015-2020), and finally to modern language-guided methods (2020-present). This section provides a comprehensive review of this evolution, establishing the foundational context necessary to understand current LLM-based approaches and their advantages over existing methods.
II-A Classical Crash Detection Methods
Early approaches to crash detection relied primarily on low-level image processing techniques and traditional machine learning methods. These foundational works established the core challenges and evaluation frameworks that continue to influence modern research.
II-A1 Motion-Based Detection
Classical motion-based approaches formed the backbone of early crash detection systems, leveraging fundamental computer vision techniques to identify anomalous vehicle behavior.
Optical Flow and Frame Differencing: Early systems used optical flow analysis to track vehicle trajectories and detect sudden changes indicative of crashes [1]. Kamijo et al. pioneered intersection monitoring using optical flow to identify collision events, achieving 85% accuracy on controlled datasets. Frame differencing techniques detected abrupt changes between consecutive frames, with Morris and Trivedi providing a comprehensive survey of trajectory-based analysis methods [2]. However, these approaches struggled with occlusions, varying lighting conditions, and camera jitter, limiting their real-world applicability.
Background Subtraction: Background subtraction methods isolated moving objects from static scenes to identify crash-related anomalies [3]. Piciarelli et al. developed trajectory-based anomaly detection using statistical models of normal traffic flow, achieving detection rates of 78% on highway surveillance data. While effective in controlled environments, these methods suffered from sensitivity to environmental changes and required extensive parameter tuning for different scenarios.
Trajectory Analysis: Trajectory-based methods analyzed vehicle paths to identify deviations from normal traffic patterns. Early work by Jiang et al. used spatiotemporal context to detect anomalous events, incorporating both spatial and temporal features for improved robustness [20]. Calderara et al. employed spectral graph analysis to model normal trajectory patterns, detecting crashes as significant deviations from learned behaviors [21]. These approaches achieved 70-80% accuracy on controlled datasets but struggled with complex multi-vehicle scenarios.
II-A2 Traditional Machine Learning Approaches
As computational power increased, researchers began applying traditional machine learning techniques to crash detection, moving beyond simple heuristics to data-driven approaches.
Support Vector Machines (SVMs): Li et al. pioneered the use of SVMs for crash injury severity classification, demonstrating superior performance compared to traditional statistical methods [22]. Their approach achieved 82% accuracy on real-world crash data, establishing SVMs as a viable tool for crash-related prediction tasks. However, SVMs required careful feature engineering and struggled with temporal dependencies in video data.
Neural Networks: Early neural network approaches, such as those developed by Abdelwahab and Abdel-Aty, used artificial neural networks to predict driver injury severity at signalized intersections [4]. These methods achieved 76% accuracy and demonstrated the potential of learning-based approaches, though they were limited by computational constraints and small datasets.
Accident Detection Systems: Ki and Lee developed comprehensive accident detection systems using image processing combined with traditional pattern recognition techniques [23]. Their system integrated multiple visual cues including motion patterns, shape analysis, and temporal consistency, achieving 83% detection accuracy on controlled test scenarios. However, the system required manual parameter tuning and failed to generalize across different traffic scenarios.
II-A3 Performance Analysis and Limitations
Classical approaches established important benchmarks and revealed fundamental challenges that continue to influence modern research.
Performance Characteristics: Classical methods typically achieved 70-85% accuracy on controlled datasets, with significant performance degradation in real-world scenarios. Motion-based approaches excelled in detecting obvious crashes with clear trajectory changes but struggled with subtle incidents or partially occluded events. Traditional ML methods showed better generalization but required extensive feature engineering and domain expertise.
Fundamental Limitations: Several critical limitations motivated the transition to deep learning approaches:
  • Feature Engineering Bottleneck: Manual feature design required extensive domain knowledge and failed to capture complex spatiotemporal patterns.
  • Environmental Sensitivity: Performance degraded significantly under varying lighting conditions, weather changes, or camera viewpoints.
  • Occlusion Handling: Most methods failed when key visual information was occluded by other vehicles or objects.
  • Temporal Modeling: Limited ability to model long-term temporal dependencies crucial for understanding crash sequences.
  • Generalization: Poor transfer performance across different geographic regions, traffic patterns, or camera setups.
These limitations directly motivated the development of deep learning approaches, which promised to learn features automatically and handle complex spatiotemporal patterns more effectively.
The transition from classical to deep learning approaches marked a paradigm shift in video analysis, moving from hand-crafted features to learned representations. This evolution was particularly impactful for crash detection, where complex spatiotemporal patterns required more sophisticated modeling approaches.
II-A4 CNN-Based Video Understanding (2014-2018)
The introduction of convolutional neural networks to video analysis revolutionized the field, providing automatic feature learning capabilities that surpassed hand-crafted approaches.
Two-Stream Networks: Simonyan and Zisserman introduced the two-stream architecture, processing RGB frames and optical flow separately before fusion [6]. This approach achieved 88% accuracy on action recognition benchmarks, demonstrating the value of explicitly modeling motion alongside appearance. For crash detection, two-stream networks enabled better understanding of both vehicle appearance and motion dynamics, though they required pre-computed optical flow.
3D Convolutional Networks: Tran et al. pioneered 3D CNNs for spatiotemporal feature learning, extending 2D convolutions to capture temporal relationships directly [7]. Their C3D architecture achieved 85% accuracy on video classification tasks and became a foundation for many crash detection systems. 3D CNNs eliminated the need for separate motion estimation while learning temporal patterns end-to-end.
Temporal Segment Networks: Wang et al. developed TSN to model long-range temporal structure by sampling sparse segments from videos [24]. This approach achieved state-of-the-art performance while being computationally efficient, making it suitable for crash detection applications requiring long temporal context. TSN demonstrated that sparse sampling could capture essential temporal information without processing every frame.
II-A5 Advanced Architectures (2017-2020)
As computational resources increased, researchers developed more sophisticated architectures specifically designed for complex video understanding tasks.
I3D and Kinetics: Carreira and Zisserman introduced Inflated 3D ConvNets (I3D), inflating successful 2D architectures into 3D while pre-training on the large-scale Kinetics dataset [8]. I3D achieved 95% accuracy on action recognition, establishing new benchmarks for video understanding. For crash detection, I3D provided robust spatiotemporal features that could capture subtle vehicle interactions and collision dynamics.
SlowFast Networks: Feichtenhofer et al. developed SlowFast networks to capture both slow semantic and fast motion information through dual-pathway architectures [9]. This approach achieved superior performance on action recognition while being computationally efficient. SlowFast networks proved particularly valuable for crash detection, where both slow semantic understanding (vehicle types, road conditions) and fast motion analysis (sudden movements, impacts) are crucial.
II-A6 Video Anomaly Detection (2018-2021)
Parallel developments in anomaly detection provided direct foundations for crash detection research, establishing evaluation protocols and baseline methods.
Real-World Anomaly Detection: Sultani et al. introduced the UCF-Crime dataset and established weakly-supervised learning for anomaly detection [10]. Their approach achieved 82% AUC on real-world surveillance videos, including traffic accidents. This work demonstrated the feasibility of learning from weakly-labeled data, crucial for crash detection where frame-level annotations are expensive.
Future Frame Prediction: Liu et al. proposed using future frame prediction for anomaly detection, learning to predict normal patterns and flagging deviations [11]. This approach achieved 84% AUC on standard benchmarks and provided a natural framework for crash anticipation. The method showed that predictive models could identify anomalies before they fully manifested.
Memory-Guided Learning: Park et al. developed memory-guided normality learning, using external memory to store prototypical normal patterns [12]. Their approach achieved 88% AUC on anomaly detection benchmarks, demonstrating the value of explicit normal pattern modeling. For crash detection, memory-guided approaches enabled better understanding of normal traffic flow and more accurate anomaly identification.
II-A7 Performance Evolution and Remaining Challenges
Deep learning approaches achieved significant performance improvements over classical methods, but revealed new challenges that motivated the transition to multimodal approaches.
Performance Gains: Deep learning methods typically achieved 85-95% accuracy on video analysis benchmarks, representing 10-20% improvements over classical approaches. CNN-based crash detection systems demonstrated superior robustness to environmental variations and better generalization across different scenarios.
Remaining Limitations: Despite significant advances, deep learning approaches faced several challenges:
  • Interpretability: Black-box nature made it difficult to understand why certain decisions were made, crucial for safety-critical applications.
  • Data Hunger: Required large annotated datasets that were expensive and time-consuming to create for crash scenarios.
  • Context Understanding: Limited ability to incorporate high-level contextual information and causal reasoning.
  • Long-term Dependencies: Difficulty modeling very long temporal sequences due to computational constraints.
These limitations motivated researchers to explore multimodal approaches that could incorporate linguistic reasoning and contextual understanding, leading to the development of vision-language models.
II-B Multimodal Learning Foundations
The emergence of multimodal learning represented a crucial bridge between pure computer vision approaches and modern language-guided systems. Understanding these foundations is essential for appreciating how current LLM-based crash detection systems achieve their capabilities.
II-B1 Theoretical Foundations
Multimodal learning theory, as comprehensively surveyed by Baltrušaitis et al., established the mathematical and conceptual frameworks for combining information from multiple modalities [25].
Representation Learning: The core challenge in multimodal learning is learning joint representations that capture complementary information across modalities. Ngiam et al. pioneered deep multimodal learning, demonstrating that shared representations could improve performance even when some modalities were missing during inference [26]. This work established that multimodal systems could be more robust than unimodal approaches, a principle crucial for crash detection systems that must operate under varying sensor availability.
Fusion Strategies: Ramachandram and Taylor provided a comprehensive taxonomy of fusion approaches, categorizing them into early fusion (feature-level), late fusion (decision-level), and hybrid approaches [27]. Early fusion combines raw features before processing, enabling deep interaction between modalities but increasing computational complexity. Late fusion processes modalities independently before combining decisions, offering modularity but potentially missing cross-modal interactions. These fusion strategies directly influence modern LLM-based crash detection architectures.
II-B2 Early Vision-Language Models (2019-2022)
The development of vision-language models established the technical foundations for incorporating linguistic reasoning into visual understanding tasks.
UNITER and Universal Representation: Chen et al. introduced UNITER, demonstrating universal image-text representation learning through careful pre-training on diverse datasets [13]. UNITER achieved state-of-the-art performance on multiple vision-language tasks, establishing the viability of unified multimodal architectures. For crash detection, UNITER’s approach showed how textual descriptions could enhance visual understanding of complex scenarios.
Align Before Fuse: Li et al. developed alignment-based approaches that explicitly aligned visual and textual representations before fusion [14]. Their momentum distillation technique achieved superior performance by ensuring semantic consistency between modalities. This work established alignment as a crucial component for multimodal systems, directly influencing how modern crash detection systems synchronize visual events with textual descriptions.
Simple Visual Language Models: Wang et al. demonstrated that simple architectures could achieve competitive performance through careful pre-training, introducing SimVLM [28]. This work showed that architectural complexity was less important than training methodology, influencing the development of efficient crash detection systems suitable for real-time deployment.
II-B3 Video-Language Understanding
The extension of vision-language models to video represented a crucial step toward current crash detection capabilities.
ClipBERT: Lei et al. introduced ClipBERT for video-language learning, demonstrating that sparse sampling could enable efficient processing of long video sequences [15]. ClipBERT achieved 89% accuracy on video question answering while being computationally efficient. For crash detection, this approach showed how to process extended temporal sequences without prohibitive computational costs.
VideoCLIP: Xu et al. developed VideoCLIP for zero-shot video-text understanding, extending CLIP’s contrastive learning approach to video [16]. VideoCLIP demonstrated strong transfer capabilities, achieving good performance on unseen video categories. This work established the potential for crash detection systems to generalize to new incident types without additional training.
Unified Video-Language Pre-training: Wang et al. explored unified approaches that jointly processed video and language during pre-training [29]. Their work demonstrated that unified pre-training could improve both video understanding and language generation, establishing foundations for systems that both detect crashes and generate explanatory descriptions.
II-B4 Transportation-Specific Datasets
The development of large-scale transportation datasets enabled the application of multimodal learning to automotive scenarios.
Cityscapes: Cordts et al. introduced Cityscapes for semantic urban scene understanding, providing high-quality annotations for urban driving scenarios [30]. While not crash-specific, Cityscapes established evaluation protocols and baseline methods for automotive vision applications.
nuScenes: Caesar et al. developed nuScenes as a comprehensive multimodal dataset for autonomous driving, including LiDAR, radar, and camera data with detailed annotations [31]. nuScenes demonstrated the value of multimodal data for understanding complex driving scenarios, though it focused on normal driving rather than crash events.
KITTI: Geiger et al. created the KITTI benchmark suite, establishing standardized evaluation protocols for automotive vision tasks [32]. KITTI’s influence extended beyond its specific tasks, establishing methodological standards that continue to influence crash detection research.
II-B5 Edge Computing and Real-time Constraints
The practical deployment of multimodal systems required addressing computational constraints, leading to important developments in edge computing for AI applications.
Edge Computing Foundations: Shi et al. established the theoretical foundations of edge computing, identifying key challenges and opportunities for deploying AI systems at the network edge [33]. This work was crucial for understanding how crash detection systems could be deployed in vehicles and roadside infrastructure.
Edge Intelligence: Zhou et al. developed comprehensive frameworks for edge intelligence, demonstrating how to balance computational efficiency with performance [34]. Their work established design principles for deploying complex AI systems in resource-constrained environments, directly relevant to real-time crash detection requirements.
II-B6 Bridging to Modern LLM Approaches
The foundations established by multimodal learning research created the necessary conditions for the emergence of LLM-based crash detection systems.
Key Enablers: Several developments were crucial: (1) Effective fusion strategies for combining visual and textual information, (2) Pre-training methodologies that could leverage large-scale multimodal datasets, (3) Efficient architectures suitable for real-time deployment, and (4) Evaluation frameworks that could assess both detection accuracy and explanation quality.
Remaining Gaps: Despite these advances, several gaps motivated the development of LLM-based approaches: (1) Limited causal reasoning capabilities, (2) Difficulty generating human-readable explanations, (3) Poor generalization to rare or novel incident types, and (4) Inability to incorporate world knowledge and contextual understanding.
These foundations set the stage for modern LLM-based crash detection systems, which build upon multimodal learning principles while adding sophisticated language understanding and generation capabilities.
@@ II-B6 Bridging to Modern LLM Approaches
Domain Fine-tunesCRASH
@@
II-C Transition to Language-Guided Approaches
The evolution from pure computer vision to language-guided approaches represents a fundamental paradigm shift in crash detection research. This transition was motivated by the limitations identified in classical and deep learning approaches, and enabled by advances in multimodal learning foundations discussed in Section II-B.
II-C1 Motivation for Language Integration
The integration of language understanding into crash detection systems addresses several critical limitations of previous approaches:
Interpretability Gap: Classical and deep learning methods provided limited explanations for their decisions, making them unsuitable for safety-critical applications where understanding the reasoning behind crash detection is crucial for legal and insurance purposes.
Contextual Understanding: Previous methods struggled to incorporate high-level contextual information such as weather conditions, traffic patterns, or unusual circumstances that could influence crash likelihood or severity.
Generalization to Rare Events: Traditional approaches required extensive training data for each type of crash scenario, making them poorly suited for detecting rare or novel incident types that were not well-represented in training datasets.
Causal Reasoning: Earlier methods could detect that a crash occurred but struggled to understand and explain the causal chain of events leading to the incident, limiting their utility for prevention and analysis.
II-C2 Early Vision-Language Integration
The first attempts to integrate language understanding with video analysis for crash detection built upon the multimodal learning foundations established in the broader computer vision community.
CLIP-based Approaches: Researchers began experimenting with CLIP-style contrastive learning for crash detection, using textual prompts like “car accident” or “normal driving” to classify video frames [46, 47]. While limited to frame-level analysis, these approaches demonstrated the potential for zero-shot generalization to new crash types through textual descriptions.
Video-Language Models: The development of video-specific language models like Video-LLaMA marked a crucial step toward temporal reasoning in crash detection [38]. These models could process video sequences while incorporating textual guidance, enabling detection of crash-related patterns through combined visual and linguistic understanding.
Alignment-based Methods: Following the success of alignment-based approaches in general multimodal learning [14], researchers developed methods to synchronize visual crash events with textual descriptions, addressing the temporal alignment challenges specific to dynamic traffic scenarios.
This transition period established the technical feasibility of language-guided crash detection while revealing the need for more sophisticated language understanding capabilities, setting the stage for the emergence of LLM-based approaches.
II-D Large Language Models for Video Understanding